OpenStudy (anonymous):

There is a circle with an 8-foot chord. The midpoint of the chord is located 3 feet from the center of the circle. What is the length of the diameter of the circle?

OpenStudy (anonymous):

Have you read the properties of circles and the Pythagoras' theorem ?

OpenStudy (anonymous):

yes math is not my subject

OpenStudy (anonymous):

you want the method straight-away ?

OpenStudy (anonymous):

i just need help

OpenStudy (anonymous):

let me draw the fig for you 

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

according to a property , the perpendicular bisector of any chord in the circle passes through the centre of the circle. So, the triangle here is a right-angled triangle. Now, can you solve ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it 8

OpenStudy (anonymous):

Now, we can apply the Pythagoras' theorem to this triangle and find out value of r. Then our diameter would be d = 2 x r

OpenStudy (anonymous):



OpenStudy (anonymous):

it's not 8. Can you write the Pythagoras' formula ?

OpenStudy (anonymous):

its 11

OpenStudy (anonymous):

The formula is : \[h^2 = p^2 + b^2\] where h= hypotenuse of the right traingle, and p, b the perpendicular and base respectively. So, here p=4, b=3 then can you find h,( or as in our case r ) ?

OpenStudy (anonymous):

Ask me if you didn't get the explanation .

OpenStudy (anonymous):

well, you're not responding. \[r = \sqrt{4^2 +3^2}\]\[ r = \sqrt{25} = 5\]So, the ans is dia = 2x5 ft =10 ft

OpenStudy (anonymous):

thanks